# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

### v1

@@ III-F Non-Parametric Cross-Guided Pooling
colsep=2pt, rowsep = 0.05cm, column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, column13 = c, column14 = c, column15 = c, row16 = bg=lightgray, row30 = bg=lightgray, cell11 = c=2,r=2l, cell13 = r=2, cell15 = c=3, cell19 = c=3, cell113 = c=3, cell31 = r=14c, bg=lightgray, cell171 = r=14c, bg=lightgray, hline1,17,31 = -0.08em, hline2 = 5-7,9-11,13-15, hline3 = -, Methods & Trainable
Param TCGA-NSCLC TCGA-BRCA TCGA-RCC
AUC F1 ACC AUC F1 ACC AUC F1 ACC
CLIP ImageNet Pretrained Max-pooling 197K 63.80±6.84 60.40±4.76 60.70±4.75 60.42±4.35 56.40±3.58 68.552±6.54 84.51±3.21 65.83±2.72 69.26±2.33
Mean-pooling 197K 69.53±4.74 63.76±5.77 64.69±4.31 66.64±2.41 60.70±2.78 71.73±3.59 93.31±0.66 78.64±0.74 81.29±1.08
ABMIL [32] 461K 66.95±4.31 62.60±3.75 62.96±3.65 67.92±3.90 61.72±3.60 72.77±3.15 93.41±1.41 79.80±1.56 82.47±1.46
CLAM-SB [31] 660K 67.49±5.94 62.86±4.19 63.51±4.19 67.80±5.14 60.51±5.07 72.46±4.36 93.85±1.52 79.87±3.17 83.21±2.67
CLAM-MB [31] 660K 69.65±3.61 64.52±3.22 65.14±2.69 67.98±4.86 60.68±6.47 74.09±3.52 93.59±1.16 78.72±2.18 8103±2.06
TransMIL [1] 2.54M 64.82±8.01 59.17±10.87 62.00±5.18 65.31±6.02 57.72±2.48 68.12±4.11 94.17±1.23 79.63±1.52 81.86±1.41
DSMIL [33] 462K 66.00±9.23 63.87±7.00 64.11±6.65 66.18±10.08 59.35±8.01 67.52±11.56 91.53±5.17 78.38±6.56 80.69±6.47
RRT-MIL [40] 2.63M 66.47±6.73 62.10±6.17 63.20±5.24 66.33±4.30 61.14±5.93 71.21±8.94 93.89±1.91 81.04±2.11 83.30±2.24
CoOp [25] 337K 69.06±4.06 63.87±3.77 64.27±3.55 68.86±3.45 61.64±2.40 72.10±3.22 94.18±1.72 79.88±2.40 82.15±1.96
CoCoOp [24] 370K 64.37±2.28 60.95±1.55 61.37±1.36 66.50±3.02 59.64±2.90 71.07±4.93 85.68±2.66 67.72±3.49 71.00±2.90
Metaprompt [26] 360K 75.94±3.01 70.35±3.09 70.41±3.09 69.12±4.12 63.39±4.28 74.65±7.20 94.18±1.56 80.03±2.06 82.52±2.15
TOP [8] 2.11M 73.56±3.14 68.19±1.22 68.77±2.53 69.75±4.66 61.32±6.12 71.68±2.56 93.56±1.22 79.66±1.97 80.79±1.05
ViLa-MIL [36] 2.77M 74.85±7.62 68.74±5.86 68.87±5.97 70.13±3.86 62.04±2.28 71.93±2.31 93.34±1.49 79.40±1.13 81.81±0.92
[…]
PLIP Pathology Pretrained Max-pooling 197K 71.78±4.13 66.40±3.51 66.66±3.42 66.66±2.36 60.32±2.24 71.57±4.83 95.18±0.63 81.63±0.92 84.30±1.30
Mean-pooling 197K 70.55±6.64 65.32±5.60 65.50±5.55 71.62±2.41 64.62±2.96 74.45±2.49 94.75±0.51 82.22±0.67 85.24±0.80
ABMIL [32] 461K 78.54±4.29 72.06±3.79 72.12±3.78 72.18±1.28 64.49±1.74 74.63±1.31 96.51±0.63 85.66±1.97 87.94±1.92
CLAM-SB [31] 660K 80.56±4.57 73.15±4.05 73.27±3.97 73.49±2.12 65.22±2.61 75.05±3.88 96.41±0.36 84.71±1.60 87.25±1.34
CLAM-MB [31] 660K 80.68±3.63 73.15±3.00 73.32±2.83 74.33±1.76 66.11±1.94 76.11±2.03 96.58±0.59 85.20±0.83 87.85±0.79
TransMIL [1] 2.54M 73.40±10.33 66.92±7.94 67.21±7.63 70.52±2.45 62.06±1.67 70.14±2.77 96.35±0.54 83.70±0.80 86.33±0.46
DSMIL [33] 462K 77.75±7.22 72.84±6.31 73.08±6.00 70.14±4.11 63.01±2.78 71.48±5.37 93.01±6.05 79.58±9.16 82.87±7.32
RRT-MIL [40] 2.63M 76.30±10.01 70.86±7.47 71.01±7.44 72.77±2.20 65.74±2.34 74.38±4.01 96.09±1.06 83.94±2.05 86.56±2.28
CoOp [25] 337K 77.92±5.48 71.58±4.74 71.63±4.75 73.77±2.83 64.88±1.26 74.14±3.38 95.76±0.80 83.23±2.07 85.90±1.63
CoCoOp [24] 370K 72.62±8.45 66.63±5.83 66.97±5.85 71.21±4.20 62.95±3.95 73.57±6.31 95.81±0.42 83.18±1.35 86.02±1.03
Metaprompt [26] 360K 78.31±5.66 72.03±4.60 71.86±4.61 73.98±2.15 65.50±2.05 75.56±4.58 95.75±0.48 83.52±1.46 86.62±1.43
TOP [8] 2.11M 78.91±3.79 72.33±4.89 72.91±4.61 74.06±2.66 65.17±2.16 76.51±1.79 95.06±0.51 82.86±1.35 86.14±0.98
ViLa-MIL [36] 2.77M 80.98±2.52 73.81±3.64 73.94±3.56 74.86±2.45 66.03±1.81 77.35±1.63 95.72±0.60 83.85±1.10 86.53±1.03
MSCPT(ours) 1.35M 84.29±3.97 76.39±5.69 76.54±5.49 75.55±5.25 67.46±2.43 79.14±2.63 96.94±0.36 87.01±1.51 89.28±1.22
IV Experimental Results
IV-A Experimental Settings
IV-A1 Datasets
To comprehensively assess the performance of our Multi-Scale and Context-focused Prompt Tuning (MSCPT), three real datasets from the Cancer Genome Atlas (TCGA) Data Portal were used: TCGA-NSCLC, TCGA-BRCA, and TCGA-RCC.
TCGA-NSCLC is a dataset of 1041 non-small cell lung cancer (NSCLC) WSIs, including 530 lung adenocarcinoma (LUAD) and 511 lung squamous cell carcinoma (LUSC) slides. 20% of the dataset (209 slides) is used for training, and the remaining 80% (832 slides) is used for testing.
TCGA-BRCA is a dataset comprising 1056 slides of breast invasive carcinoma (BRCA) WSIs. This dataset includes 845 slides of invasive ductal carcinoma (IDC) and 211 slides of invasive lobular carcinoma (ILC). 20% of them (223 slides) are randomly selected as the training set, and the remaining 80% (833 slides) are used as the testing set.
TCGA-RCC is a renal cell carcinoma (RCC) WSIs dataset containing 873 slides. Precisely, it consists of 121 slides of chromophobe renal cell carcinoma (CHRCC), 455 slides of clear-cell renal cell carcinoma (CCRCC), and 297 slides of papillary renal cell carcinoma (PRCC). Likewise, 20% of the dataset (175 slides) is randomly taken out for training, while 698 slides are reserved for testing.
IV-A2 Evaluation Metrics
@@
For all datasets, we leverage Accuracy (ACC), Area Under Curve (AUC), and macro F1-score (F1) to evaluate model performance. To reduce the impact of data split on model evaluation, we follow ViLa-MIL [36] and employ five fixed seeds to perform five rounds of dataset splitting, model training, and testing. We report the mean and standard deviation of the metrics over five seeds.
IV-A3 Model Zoo

### v2

@@ 3.6 Non-Parametric Cross-Guided Pooling
colsep=1.5pt, rowsep = 0.05cm, column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, column13 = c, column14 = c, column15 = c, column16 = c, column17 = c, column18 = c, column19 = c, column20 = c, column21 = c, column22 = c, column23 = c, row16 = bg=lightgray, row30 = bg=lightgray, row44 = bg=lightgray, cell11 = c=2,r=2c, cell13 = r=2, cell14 = c=3, cell18 = c=3, cell112 = c=3, cell116 = c=3, cell120 = c=3, cell31 = r=14c, bg=lightgray, cell171 = r=14c, bg=lightgray, cell311 = r=14c, bg=lightgray, hline1,17,31,45 = -0.08em, hline2 = 4-6,8-10,12-14,16-18,20-22, hline3 = -, Methods & Train. Param TCGA-NSCLC
(Subtyping, 2 classes) TCGA-RCC
(Subtyping, 3 classes) UBC-OCEAN
(Subtyping, 5 classes) TCGA-BRCA
(Recurrence Prediction, 2 classes) PANDA
(Grading, 3 classes)
AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC
CLIP Pretrained on 400M Natural Image Pairs Max-pooling 197K 63.80±6.84 60.40±4.76 60.70±4.75 84.51±3.21 65.83±2.72 69.26±2.33 87.76±1.45 62.36±5.63 63.58±4.05 59.74±1.88 57.03±1.71 63.85±2.06 72.12±1.71 51.68±3.55 56.48±1.50
Mean-pooling 197K 69.53±4.74 63.76±5.77 64.69±4.31 93.31±0.66 78.64±0.74 81.29±1.08 85.18±1.70 59.59±4.02 62.16±3.75 64.19±3.40 58.08±3.38 61.01±4.19 77.85±1.32 58.86±1.69 62.75±1.96
ABMIL[29] 461K 66.95±4.31 62.60±3.75 62.96±3.65 93.41±1.41 79.80±1.56 82.47±1.46 88.03±0.98 65.33±1.57 65.98±2.02 62.51±4.10 57.44±3.51 63.42±3.00 73.89±2.40 51.07±2.87 57.93±2.99
CLAM-SB[28] 660K 67.49±5.94 62.86±4.19 63.51±4.19 93.85±1.52 79.87±3.17 83.21±2.67 88.10±1.55 66.11±2.88 66.76±2.97 65.79±2.89 59.94±3.09 63.53±4.64 76.66±1.53 56.92±2.87 58.76±3.71
CLAM-MB[28] 660K 69.65±3.61 64.52±3.22 65.14±2.69 93.59±1.16 78.72±2.18 81.03±2.06 87.09±0.73 63.54±2.20 65.69±1.64 64.74±3.37 59.59±3.23 64.12±4.83 74.68±2.43 55.40±3.06 58.08±3.61
TransMIL[1] 2.54M 64.82±8.01 59.17±10.87 62.00±5.18 94.17±1.23 79.63±1.52 81.86±1.41 83.49±1.77 54.36±4.74 55.59±5.88 60.85±4.37 55.97±3.21 60.93±3.39 71.95±1.86 53.68±3.00 57.43±3.23
DSMIL[30] 462K 66.00±9.23 63.87±7.00 64.11±6.65 91.53±5.17 78.38±6.56 80.69±6.47 87.55±0.55 64.84±1.80 66.23±1.89 60.42±3.07 57.82±2.58 63.44±2.52 72.14±3.36 51.79±6.61 58.69±4.25
RRT-MIL[40] 2.63M 66.47±6.73 62.10±6.17 63.20±5.24 93.89±1.91 81.04±2.11 83.30±2.24 88.02±1.10 62.40±2.79 64.22±3.02 62.38±4.82 57.80±5.76 61.21±4.60 74.03±5.59 53.78±7.21 60.27±5.31
CoOp[22] 337K 69.06±4.06 63.87±3.77 64.27±3.55 94.18±1.72 79.88±2.40 82.15±1.96 88.65±0.92 66.82±2.53 68.04±2.11 64.16±3.37 58.69±3.08 62.89±2.67 78.71±3.03 60.38±4.01 64.37±3.31
CoCoOp[21] 370K 64.37±2.28 60.95±1.55 61.37±1.36 85.68±2.66 67.72±3.49 71.00±2.90 88.57±1.07 66.93±1.69 67.70±1.70 63.62±3.79 58.76±2.39 62.78±3.55 77.01±1.81 55.02±5.33 60.56±2.48
Metaprompt[23] 688K 75.94±3.01 70.35±3.09 70.41±3.09 94.18±1.56 80.03±2.06 82.52±2.15 88.79±0.76 67.43±1.82 68.56±2.22 65.78±2.73 60.21±3.49 64.22±2.21 78.87±2.51 59.90±3.81 62.71±3.71
TOP[6] 2.11M 73.56±3.14 68.19±1.22 68.77±2.53 93.56±1.22 79.66±1.97 80.79±1.05 87.93±0.88 66.74±2.44 67.15±1.82 63.18±2.44 57.96±4.14 63.91±3.12 78.55±2.34 58.83±3.67 61.80±3.51
ViLa-MIL[33] 2.77M 74.85±7.62 68.74±5.86 68.87±5.97 93.34±1.49 79.40±1.13 81.81±0.92 87.19±1.03 64.43±1.84 65.49±1.36 64.00±2.18 57.59±3.41 63.67±1.98 79.63±1.86 61.11±3.25 65.57±2.49
MSCPT(ours) 1.35M 78.67±3.93 72.47±3.13 72.67±2.96 95.04±1.31 83.78±2.19 85.62±2.14 89.99±0.93 68.14±3.65 70.10±3.61 67.41±5.35 61.88±3.48 66.00±2.70 81.81±2.01 63.06±2.93 68.02±2.96
PLIP Pretrained on 208K Pathology Image Pairs Max-pooling 197K 71.78±4.13 66.40±3.51 66.66±3.42 95.18±0.63 81.63±0.92 84.30±1.30 91.57±1.35 71.06±3.30 71.09±3.39 65.83±2.66 59.92±1.62 66.12±3.57 78.72±3.05 59.24±3.22 62.86±3.13
Mean-pooling 197K 70.55±6.64 65.32±5.60 65.50±5.55 94.75±0.51 82.22±0.67 85.24±0.80 87.98±1.06 62.61±1.64 65.05±1.82 67.20±3.34 61.64±2.86 62.33±3.87 79.21±1.74 60.94±0.80 64.28±1.55
ABMIL[29] 461K 78.54±4.29 72.06±3.79 72.12±3.78 96.51±0.63 85.66±1.97 87.94±1.92 90.93±0.33 70.82±1.52 70.82±1.98 67.89±3.17 63.24±1.71 66.94±3.75 81.19±4.25 64.04±4.56 66.67±4.08
CLAM-SB[28] 660K 80.56±4.57 73.15±4.05 73.27±3.97 96.41±0.36 84.71±1.60 87.25±1.34 91.54±1.16 71.39±3.13 71.65±2.86 69.54±3.37 64.81±2.40 68.51±3.06 82.84±3.59 65.39±3.86 67.76±3.22
CLAM-MB[28] 660K 80.68±3.63 73.15±3.00 73.32±2.83 96.58±0.59 85.20±0.83 87.85±0.79 91.66±0.95 71.55±3.18 71.62±3.19 68.84±4.10 64.31±1.09 67.97±3.00 82.07±3.45 64.81±3.96 67.53±4.00
TransMIL[1] 2.54M 73.40±10.33 66.92±7.94 67.21±7.63 96.35±0.54 83.70±0.80 86.33±0.46 91.64±1.76 66.99±2.91 69.51±2.31 67.10±5.76 60.28±5.45 65.60±4.73 79.86±2.18 61.87±1.53 65.25±1.81
DSMIL[30] 462K 77.75±7.22 72.84±6.31 73.08±6.00 93.01±6.05 79.58±9.16 82.87±7.32 90.40±1.24 68.54±5.13 70.05±3.44 63.59±3.53 61.22±3.38 65.16±3.68 72.14±3.36 51.79±6.61 58.69±4.25
RRT-MIL[40] 2.63M 76.30±10.01 70.86±7.47 71.01±7.44 96.09±1.06 83.94±2.05 86.56±2.28 91.65±1.13 68.37±2.36 69.80±2.44 66.20±5.57 62.09±4.99 65.19±5.09 82.05±3.24 63.43±3.13 66.22±3.20
CoOp[22] 337K 77.92±5.48 71.58±4.74 71.63±4.75 95.76±0.80 83.23±2.07 85.90±1.63 91.53±0.73 71.21±3.51 71.60±3.93 68.47±4.96 64.26±2.45 66.71±3.76 81.74±2.86 64.81±2.90 67.80±3.11
CoCoOp[21] 370K 72.62±8.45 66.63±5.83 66.97±5.85 95.81±0.42 83.18±1.35 86.02±1.03 91.65±0.70 71.48±2.25 72.01±2.15 67.49±4.76 64.57±2.19 67.24±3.66 81.36±2.23 62.42±2.08 65.59±2.93
Metaprompt[23] 688K 78.31±5.66 72.03±4.60 71.86±4.61 95.75±0.48 83.52±1.46 86.62±1.43 91.44±0.30 70.90±1.95 71.12±1.85 68.77±4.54 63.97±3.38 67.85±3.18 80.94±2.32 62.19±3.48 65.70±2.96
TOP[6] 2.11M 78.91±3.79 72.33±4.89 72.91±4.61 95.06±0.51 82.86±1.35 86.14±0.98 91.18±0.76 71.16±1.24 71.67±2.08 68.52±4.07 64.73±2.65 67.70±4.01 81.03±2.49 62.87±3.57 66.46±3.15
ViLa-MIL[33] 2.77M 80.98±2.52 73.81±3.64 73.94±3.56 95.72±0.60 83.85±1.10 86.53±1.03 91.00±0.39 70.42±1.02 70.98±1.17 69.24±2.60 63.80±1.67 67.92±1.69 82.17±2.73 64.35±3.15 66.70±3.56
MSCPT(ours) 1.35M 84.29±3.97 76.39±5.69 76.54±5.49 96.94±0.36 87.01±1.51 89.28±1.22 92.59±1.92 73.10±2.71 73.65±3.67 71.49±2.68 66.99±3.78 70.19±3.11 84.90±3.85 68.39±3.96 70.55±4.50
CONCH Pretrained on 1.17M Pathology Image Pairs Max-pooling 131K 94.68±1.91 87.97±1.85 87.98±1.85 97.42±0.25 90.20±0.91 91.35±0.65 95.18±0.56 80.78±1.77 80.99±1.87 76.00±2.11 67.18±2.93 71.38±3.17 84.42±2.57 65.90±3.02 69.85±3.62
Mean-pooling 131K 90.09±2.74 82.11±3.98 82.12±3.97 97.28±0.23 88.05±0.28 89.86±0.59 94.59±0.47 76.98±1.46 78.48±1.19 77.71±3.05 68.60±2.72 72.13±3.05 82.10±1.20 64.26±0.74 68.75±1.29
ABMIL[29] 329K 94.69±2.17 87.80±2.61 87.84±2.57 98.41±0.50 91.41±1.15 92.69±1.05 96.02±0.36 81.24±1.09 82.35±1.09 79.40±1.17 70.19±2.26 74.95±2.97 86.98±1.30 69.99±1.57 72.97±1.40
CLAM-SB[28] 528K 96.64±1.44 88.92±2.27 88.93±2.27 98.02±0.37 91.07±1.34 91.85±1.08 96.23±0.35 82.19±1.41 82.74±0.94 79.34±2.44 70.02±2.88 73.70±3.10 87.95±2.02 71.75±1.68 74.68±2.32
CLAM-MB[28] 528K 95.85±1.28 88.78±1.45 88.78±1.44 97.88±0.43 90.74±1.35 91.59±1.11 95.54±0.23 81.21±1.40 82.07±0.94 78.93±3.05 70.06±2.57 74.45±1.97 87.22±1.43 71.08±1.93 73.93±2.16
TransMIL[1] 2.40M 93.29±3.46 86.42±4.03 86.44±4.05 98.29±0.25 90.56±0.71 92.12±0.75 95.67±0.60 78.78±0.94 80.20±1.75 77.94±3.54 67.94±4.32 70.85±4.50 86.09±1.55 69.83±2.54 72.97±2.98
DSMIL[30] 331K 85.85±5.11 81.17±4.82 81.30±4.63 97.96±0.98 90.50±2.97 92.06±2.26 95.51±0.55 80.61±3.40 81.72±2.79 76.37±2.01 68.38±3.01 72.65±2.68 83.29±4.08 65.79±5.54 70.24±5.98
RRT-MIL[40] 2.49M 94.60±2.37 88.50±3.13 88.53±3.11 98.35±0.38 90.99±1.17 92.41±0.75 95.58±0.26 79.23±3.26 80.25±2.95 78.77±2.71 69.22±3.24 73.58±3.73 84.79±3.54 68.02±2.73 70.98±3.70
CoOp[22] 341K 92.52±2.53 84.97±2.79 85.00±2.76 98.03±0.46 90.75±0.66 92.23±0.64 95.05±0.92 78.58±1.80 80.59±1.78 78.49±2.92 69.88±2.79 74.02±2.74 86.82±1.62 70.05±1.88 73.82±1.18
CoCoOp[21] 383K 93.63±2.26 86.73±2.52 86.75±2.50 97.92±0.59 90.42±1.05 91.98±1.00 94.95±0.59 78.92±3.20 80.78±2.70 78.22±2.90 69.61±2.40 74.11±2.52 87.06±1.02 69.51±2.03 73.15±1.01
Metaprompt[23] 694K 94.98±1.16 88.05±0.94 88.08±0.92 98.28±0.39 91.62±0.98 92.51±0.87 95.84±0.55 80.59±0.74 82.11±0.83 79.07±3.22 70.06±3.64 73.65±4.25 87.59±0.48 71.72±0.44 74.61±0.35
TOP[6] 2.76M 93.77±1.25 87.96±2.69 88.11±2.86 98.15±0.34 90.79±1.53 92.18±2.13 94.97±1.14 78.68±2.89 80.30±2.94 78.88±3.37 68.70±3.94 72.91±4.72 84.51±2.11 67.76±2.85 70.03±3.33
ViLa-MIL[33] 2.32M 94.73±1.76 87.58±2.29 87.60±2.30 98.00±0.50 88.97±1.93 90.54±2.56 95.10±1.03 79.90±1.79 81.13±1.79 78.31±2.71 69.80±2.98 74.34±1.49 85.22±2.20 69.04±2.81 71.68±3.41
MSCPT(ours) 1.69M 96.45±0.94 89.68±2.71 89.71±2.70 98.66±0.29 92.50±1.18 93.45±1.16 96.14±0.79 83.05±2.04 83.58±1.94 80.81±3.10 71.71±2.79 75.86±2.41 89.39±1.71 73.58±2.84 76.66±2.61
4 Experimental Results
4.1 Experimental Settings
4.1.1 Datasets
To comprehensively evaluate MSCPT’s performance, we used five real-world datasets across three downstream tasks, collected from multiple centers. Additionally, to assess the generalizability of MSCPT across different VLMs, we conducted extensive experiments using three selected VLMs.
TCGA-NSCLS, TCGA-RCC, and UBC-OCEAN datasets are used to examine the cancer subtyping ability of MSCPT. TCGA-NSCLC contains 1041 non-small cell lung cancer WSIs, comprising 530 lung adenocarcinoma (LUAD) and 511 lung squamous cell carcinoma (LUSC) WSIs. TCGA-RCC includes 873 renal cell carcinoma WSIs: 121 chromophobe renal cell carcinoma (CHRCC), 455 clear-cell renal cell carcinoma (CCRCC), and 297 papillary renal cell carcinoma (PRCC). UBC-OCEAN comprises 538 ovarian WSIs, with 99 clear cell carcinoma (CC), 124 endometrioid carcinoma (EC), 221 high-grade serous carcinoma (HGSC), 47 low-grade serous carcinoma (LGSC), and 46 mucinous carcinoma (MC) WSIs.
The TCGA-BRCA dataset is used for predicting the risk of cancer recurrence and comprises 1,099 WSIs. Typically, the recurrence risk of breast cancer is predicted by gene expression data combined with OncotypeDx (ODX). Although the original dataset does not provide labels related to recurrence risk, Howard et al. [41] calculated the corresponding ODX scores based on the mRNA expression data provided by TCGA and classified the patients into high recurrence risk (316 WSIs) and low recurrence risk (783 WSIs).
The PANDA dataset includes 10,616 WSIs used for prostate cancer grading. Each slide is labeled as normal or assigned an International Society of Urological Pathology (ISUP) grade, resulting in six classes. Considering the difficulty of the few-shot weakly supervised WSI classification task, we introduced coarse-grained labels based on European Association of Urology (EAU) prostate cancer guidelines [42], mapping ISUP grades to risk categories: normal or grade 1 as low risk (5,558 WSIs), grade 2 and grade 3 as medium risk (2,585 WSIs), and grades 4–5 as high risk (2,473 WSIs).
4.1.2 Evaluation Metrics
@@
For all datasets, we leverage Accuracy (ACC), Area Under Curve (AUC), and macro F1-score (F1) to evaluate model performance. To reduce the impact of data split on model evaluation, we employ five fixed seeds to perform five rounds of dataset splitting, model training, and testing. We report the mean and standard deviation of the metrics.
4.1.3 Model Zoo